Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 13,750 $ 17,250 $ 31,000
Estimated variable manufacturing overhead per machine-hour $ 2.90 $ 3.70
Job P Job Q
Direct materials $ 28,000 $ 15,500
Direct labor cost $ 33,000 $ 13,500
Actual machine-hours used:
Molding 3,200 2,300
Fabrication 2,100 2,400
Total 5,300 4,700

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1–8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9–15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

1a. How much manufacturing overhead was applied from the Molding Department to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

1b. How much manufacturing overhead was applied from the Fabrication Department to Job P and how much was applied to Job Q?

1c.12. If Job P included 20 units, what was its unit product cost

1d. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

1.e14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

2.What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)
Answer #1
Predetermined overhead rates:
Molding Department 8.40 =2.9+(13750/2500)
Fabrication Department 15.20 =3.7+(17250/1500)
1a
Manufacturing overhead applied:
Job P 26880 =3200*8.4
Job Q 19320 =2300*8.4
1b
Manufacturing overhead applied:
Job P 31920 =2100*15.2
Job Q 36480 =2400*15.2
1c
Direct materials 28000
Direct labor cost 33000
Manufacturing overhead applied 58800 =26880+31920
Total manufacturing cost Job P 119800
Divide by units 20
Unit product cost Job P 5990
1d
Direct materials 15500
Direct labor cost 13500
Manufacturing overhead applied 55800 =19320+36480
Total manufacturing cost Job Q 84800
Divide by units 30
Unit product cost Job Q 2827
1e
Job P Job Q
Total manufacturing cost 119800 84800
Add: Markup @ 80% 95840 67840
Total price for the job 215640 152640
Divide by units 20 30
Selling price per unit 10782 5088
2
Cost of goods sold for March 204600 =119800+84800